EVIDENCE FOR RADIATIONAL TRANSITIONS IN THE TRIPLET MANIFOLD OF SOME AROMATIC MOLECULES

\underline{HAIFENG XU}, PHILIP M. JOHNSON, Chemistry Department, Stony Brook University, Stony Brook, NY 11974; TREVOR J. SEARS^{a}, Department of Chemistry, Brookhaven National Laboratory, Upton, NY 11973-5000.

Pump-probe delayed ionization studies on benzonitrile and phenylacetylene in a supersonic beam have shown evidence of very long-lived excited states lasting more than hundreds of microseconds. Measurements of the ionization thresholds of these states are consistent with their being low vibrational levels of the lowest triplet states. Excitation of the molecules was done with a frequency-doubled, Fourier transform-limited pulse-amplified c.w. laser, and the rotationally resolved structure of the S_{1}−S_{0} transition ensures that the excited molecules are monomers. Lacking a dissociative mechanism for energy loss in the triplet manifold of the isolated molecules, it is inferred that the lower triplet states are being populated by photon emission, possibly between triplet electronic states.

aWork at Stony Brook University was carried out under Contract No. DE-FG02-86ER13590 and at Brookhaven National Laboratory under Contract No. DE-AC02-98CH10886, both with the U.S. Department of Energy and supported by its Division of Chemical Sciences, Office of Basic Energy Sciences.